# Lower Bounds for Faithful Linear Representations of Subgroups of the Mapping Class Group

### v1

@@ 1. Introduction
The lower bound of n1n-1 from Theorem 4 is well known for the full braid group $B_{n}$. Dyer–Formanek–Grossman showed in [8, Proposition 2] that if $\rho:B_{n}\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ is faithful then one of its irreducible subquotients must also be faithful. The irreducible representations of $B_{n}$ of dimension at most n1n-1 were classified by Formanek [10]. In particular, no irreducible representation of dimension <n1<n-1 is faithful for $n\geqslant 3$.
@@
To the best of the authors knowledge, the value of $d(\Sigma_{g,1})$ is unknown even for $g=2$. The group $\operatorname{Mod}(\Sigma_{2}^{1})$ contains a natural copy of $B_{5}$, so that $d(\Sigma_{2}^{1})\geqslant d(B_{5})$. The smallest faithful representation of $B_{5}$ known in the literature seems to be the so-called Lawrence representation $B_{5}\hookrightarrow\operatorname{GL}_{10}(\mathbb{Z}[q^{\pm 1},t^{\pm 1}])$. If one assumed $d(E)\geqslant 10$ for all other cyclic extensions $E$ of $\operatorname{Mod}(\Sigma_{2,1})$ then Theorem 5 would say $d(\Sigma_{g}^{1})\geqslant 5g$, thus improving our lower bounds of 4g34g-3.
@@
Given $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ with $d$ small enough, we show that, unless $\rho$ annihilates $\operatorname{SIP}_{0}(\Sigma)$, the matrices $M_{i}=\rho(T_{a_{i}})-1$ and $N_{j}=\rho(T_{b_{j}})-1$ satisfy the relations
@@ Acknowledgments
I would like to express my deep gratitude to my advisors Julien Marché and Maxime Wolff for taking me as their student and suggesting me this problem, as well as for the incredibly helpful discussions we shared in the past year.
I should also thank Francesco Costantino for suggesting me to consider representations of the Torelli group. I am thankful to Bram Petri for communicating me the Kaufmann–Salter–Zhang–Zhong preprint. Finally, I am thankful to Gwenaël Massuyeau for communicating me Childers’ results on the groups $\operatorname{SIP}(\Sigma_{g})$ and $\operatorname{SIP}(\Sigma_{g}^{1})$.
@@ 2.1. Curves & Dehn Twists
    Free subgroups. Given $a,b\subseteq\Sigma$ with $|a\pitchfork b|\geqslant 2$, $T_{a}$ and $T_{b}$ generate a rank-$2$ free group in $\operatorname{PMod}(\Sigma)$ [9, Theorem 3.14].
It turns out free subgroups are ubiquitous in $\operatorname{PMod}(\Sigma)$. This is because of the so-called Tits-alternative: a subgroup $G\leqslant\operatorname{PMod}(\Sigma)$ is either virtually Abelian or it contains free groups [28, Theorem A]. In particular, up to taking powers, the subgroup generated by two mapping classes is almost always a free group.
@@
The idea is to show that any non-central $N\triangleleft\operatorname{Mod}(\Sigma_{g}^{b})$ contains pseudo-Anosov maps [^1] See [9, §13.2.3] for a definition., which is known since the mid 1980’s – see [26, Lemma 2.5]. This is also a consequence of the existence of all pseudo-Anosov normal subgroups of $\operatorname{Mod}(\Sigma_{g}^{b})$, a fact first established by Dahmani–Guirardel–Osin [6, Theorem 2.31].
@@ 2.2. The Torelli group & its Subgroups
The subgroup $\operatorname{SIP}(\Sigma)\leqslant\mathcal{I}(\Sigma)$ generated by simple intersection maps remains less well understood than $\mathcal{I}(\Sigma)$ and $\mathcal{K}(\Sigma)$. This is a normal subgroup of $\mathcal{I}(\Sigma)$: given a simple intersection pair $(a,b)$ and $f\in\operatorname{Mod}(\Sigma)$,
@@
When $\Sigma=\Sigma_{g}^{1}$, the images of simple intersection maps under the Johnson homomorphism were computed independently by Putman [33], Church [5] and Childers [4, Main Result 1]. Childers also computed the image of $\operatorname{SIP}(\Sigma_{g}^{1})$ under the so-called Birman–Craggs–Johnson homomorphism [4, Main Result 4].
@@ 3. Commutation Relations in the Mapping Class Group
The latter is a Gromov-hyperbolic graph on which $\operatorname{PMod}(\Sigma)$ acts by isometries [27]. Given the disjointness relations from §2.1, $\mathcal{C}(\Sigma)$ and $\Gamma(\operatorname{PMod}(\Sigma))$ are related by means of the embedding $\mathcal{C}(\Sigma)\hookrightarrow\operatorname{PMod}(\Sigma)$ taking $a\subseteq\Sigma$ to its Dehn twist $T_{a}$.
@@
Now each $W_{i+1}/W_{i}$ may be decomposed as a tensor product $W_{i+1}/W_{i}=W_{i,1}\otimes\cdots\otimes W_{i,n}$, where $\rho_{ij}:H_{j}\to\operatorname{GL}(W_{ij})$ is an irreducible representation and $\rho_{i}(h_{1},\ldots,h_{n})=\rho_{i,1}(h_{1})\otimes\cdots\otimes\rho_{i,n}(h_{n})$ – see, for example, [23, Proposition 2.3.23]. In particular, if $\dim W_{ij}=1$ then $\rho_{i}(1,\ldots,1,h_{j},1,\ldots,1)=1\otimes\cdots 1\otimes\rho_{ij}(h_{j})\otimes 1\otimes\cdots\otimes 1$ is a scalar operator for all $h_{j}\in H_{j}$.
Assume we can find $j\leqslant n$ with $\dim W_{ij}=1$ for all $i$. Then the matrix $\rho(\pi(1,\ldots,1,h_{j},1,\ldots,1))$ is upper triangular with respect to the basis from (3.3) for all $h_{j}\in H_{j}$. Since the group of upper triangular matrices is solvable, it follows $\pi(H_{i}^{(k)})\leqslant\ker\rho$ for some $k$. But $H_{i}$ is non-solvable and so $\pi(H_{i}^{(k)})\cong H_{i}^{(k)}\neq 1$, contradicting the assumption that $\rho$ is faithful.
This means that, for each $j\leqslant n$ we can find $i$ such that $\dim W_{ij}\geqslant 2$. It follows
$\begin{split}d&=\dim W_{1}/W_{0}+\cdots+\dim W_{p}/W_{p-1}\\ &\geqslant 2^{\#\{j:\dim W_{1,j}\geqslant 2\}}+\cdots+2^{\#\{j:\dim W_{p,j}\geqslant 2\}}\\ &\geqslant 2\cdot\#\{j:\dim W_{1,j}\geqslant 2\}+\cdots+2\cdot\#\{j:\dim W_{p,j}\geqslant 2\}\\ &\geqslant 2n,\end{split}$
as desired. ∎
@@
Figure 4. The curves $a_{1},\ldots,a_{k},b_{1},\ldots,b_{k}\subseteq\Sigma_{g}^{1}$.
In that case, consider $a_{1},\ldots,a_{k},b_{1},\ldots,b_{k}\subseteq\Sigma_{0,n}$ as in Figure 4. It follows from the discussion above that the subgroup generated by $T_{a_{1}},\ldots,T_{a_{k}},T_{b_{1}},\ldots,T_{b_{k}}\in\operatorname{PMod}(\mathbb{D}_{n})$ is a quotient of $G_{0,k}$, the direct product of $k$ copies of $F_{2}$. What is more, the projection of each $F_{2}$-factor onto its image in $\operatorname{PMod}(\Sigma_{0,n}^{1})$ is an isomorphism. Proposition 3.2 then says $d\geqslant 2k=n-1$, as desired. ∎
@@
Let $\rho:\operatorname{Mod}(\Sigma_{g}^{1})\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ be a faithful representation and $g^{\prime}=\lfloor\nicefrac{{g}}{{n}}\rfloor$. Take $d_{\operatorname{min}}=\min\{d(E):E\text{ is a cyclic extension of }\operatorname{Mod}(\Sigma_{g^{\prime},1})\}$. We want to show $d\geqslant n\cdot d_{\operatorname{min}}$. By passing to a smaller subsurface $\Sigma_{n\cdot g^{\prime}}^{1}\subseteq\Sigma_{g}^{1}$ if necessary, we may assume $g=n\cdot g^{\prime}$ with $g^{\prime}\geqslant 2$. In that case, we may view $\Sigma_{g}^{1}$ as an $(n+1)$-hole sphere with $n$ copies $S_{1},\ldots,S_{n}$ of $\Sigma_{g^{\prime}}^{1}$ attached along their boundaries, as in Figure 5.
$S_{1}$$S_{2}$$S_{n}$
Figure 5. The subsurfaces $S_{1},\ldots,S_{n}\subseteq\Sigma_{g}^{1}$.
@@
so that the action $\rho_{i}:\operatorname{Mod}(S_{1})\times\cdots\times\operatorname{Mod}(S_{n})\to\operatorname{GL}(W_{i+1}/W_{i})$ of $\operatorname{Mod}(S_{1})\times\cdots\times\operatorname{Mod}(S_{n})$ on each successive quotient is irreducible. Set $W_{i+1}/W_{i}=W_{i,1}\otimes\cdots\otimes W_{i,n}$, where $\rho_{ij}:\operatorname{Mod}(S_{j})\to\operatorname{GL}(W_{ij})$ is an irreducible representation and $\rho_{i}(f_{1},\ldots,f_{n})=\rho_{i,1}(f_{1})\otimes\cdots\otimes\rho_{i,n}(f_{n})$ [23, Proposition 2.3.23].
@@
$\begin{split}d&=\dim W_{1}/W_{0}+\cdots+\dim W_{p}/W_{p-1}\\ &\geqslant d_{\operatorname{min}}^{\#\{j:\dim W_{1,j}\geqslant d_{\operatorname{min}}\}}+\cdots+d_{\operatorname{min}}^{\#\{j:\dim W_{p,j}\geqslant d_{\operatorname{min}}\}}\\ &\geqslant d_{\operatorname{min}}\cdot\#\{(i,j):\dim W_{i,j}\geqslant d_{\operatorname{min}}\}\\ &\geqslant n\cdot d_{\operatorname{min}},\end{split}$
@@
Take $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ with $d$ small enough. In §4 and §5 we will show that, unless $\rho$ annihilates the generators of $\operatorname{SIP}_{0}(\Sigma)$, the matrices $M_{i}=\rho(T_{a_{i}})-1$ and $N_{j}=\rho(T_{b_{j}})-1$ satisfy the relations
@@
Let $M_{1},\ldots,M_{n},N_{1},\ldots,N_{n}\in M_{d}(\mathbb{C})$ be nonzero operators subject to relations (3.5). Then $\dim\sum_{i}(\operatorname{range}M_{i}+\operatorname{range}N_{i})\geqslant 3n-d$. In particular, $2d\geqslant 3n$.
@@ 4. Eigenspaces of $T_{a}$
Let $\Sigma$ be a genus $g$ surface and $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$. Take $a,b\subseteq\Sigma$ disjoint with $a$ nonseparating satisfying either (1) or (2) from Proposition 4.2. Then $(L_{a}-1)(L_{b}-1)=0$.
@@
Now since $\rho$ is nontrivial, $\dim E_{1}^{a}>2g-2$ by Proposition 4.2 and thus $\dim\mathbb{C}^{d}/E^{a}_{1}<2g-2$. It follows from Theorem 2.7 that $\bar{\rho}(f)=1$. Given $b\subseteq\Sigma_{a}$ nonseparating, we may thus write
@@
On the one hand, $\operatorname{Ext}^{1}_{\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]}(H_{\mathbb{Z}},H_{\mathbb{Z}})=H^{1}(\operatorname{Mod}(\Sigma^{\prime});\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},H_{\mathbb{Z}}))$ [3, Proposition 2.2]. Here $\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]$ acts on $\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},H_{\mathbb{Z}})$ by $f\cdot A=\Psi(f)\circ A\circ\Psi(f)^{-1}$ for all $f\in\operatorname{Mod}(\Sigma^{\prime})$ and $A\in\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},H_{\mathbb{Z}})$. Hence $\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},H_{\mathbb{Z}})\cong H_{\mathbb{Z}}^{*}\otimes_{\mathbb{Z}}H_{\mathbb{Z}}$ as $\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]$-modules, where $\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]$ acts on $H_{\mathbb{Z}}^{*}=\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},\mathbb{Z})$ via $f\cdot A=A\circ\Psi(f)^{-1}$.

### v2

@@ 1. Introduction
The lower bound of n1n-1 from Theorem 4 is well known for the full braid group $B_{n}$. Dyer–Formanek–Grossman showed in [9, Proposition 2] that if $\rho:B_{n}\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ is faithful then one of its irreducible subquotients must also be faithful. The irreducible representations of $B_{n}$ of dimension at most n1n-1 were classified by Formanek [11]. In particular, no irreducible representation of dimension <n1<n-1 is faithful for $n\geqslant 3$.
@@
To the best of the authors knowledge, the value of $d(\Sigma_{g,1})$ is unknown even for $g=2$. The group $\operatorname{Mod}(\Sigma_{2}^{1})$ contains a natural copy of $B_{5}$, so that $d(\Sigma_{2}^{1})\geqslant d(B_{5})$. The smallest faithful representation of $B_{5}$ known in the literature seems to be the so-called Lawrence representation $B_{5}\hookrightarrow\operatorname{GL}_{10}(\mathbb{Z}[q^{\pm 1},t^{\pm 1}])$ [1, 25]. If one assumed $d(E)\geqslant 10$ for all other cyclic extensions $E$ of $\operatorname{Mod}(\Sigma_{2,1})$ then Theorem 5 would say $d(\Sigma_{g}^{1})\geqslant 5g$, thus improving our lower bounds of 4g34g-3.
@@
Given $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ with $d$ small enough, we show that, unless $\operatorname{SIP}_{0}(\Sigma)\leqslant\ker\rho$, the matrices $M_{i}=\rho(T_{a_{i}})-1$ and $N_{j}=\rho(T_{b_{j}})-1$ satisfy the “annihilation relations”
@@ Acknowledgments
I would like to express my deep gratitude to my advisors J. Marché and M. Wolff for taking me as their student and suggesting me this problem, as well as for the incredibly helpful discussions we shared in the past year.
I should also thank F. Costantino for suggesting me to consider representations of the Torelli group. I am thankful to B. Petri for communicating me the Kaufmann–Salter–Zhang–Zhong preprint, and to G. Massuyeau for communicating me Childers’ results on the groups $\operatorname{SIP}(\Sigma_{g})$ and $\operatorname{SIP}(\Sigma_{g}^{1})$. Finally, I am grateful to J. Martel and the anonymous referees for their helpful suggestions on earlier drafts.
@@ 2.1. Curves & Dehn Twists
    Free subgroups. Given $a,b\subseteq\Sigma$ with $|a\pitchfork b|\geqslant 2$, $T_{a}$ and $T_{b}$ generate a rank-$2$ free group in $\operatorname{PMod}(\Sigma)$ [10, Theorem 3.14].
It turns out free subgroups are ubiquitous in $\operatorname{PMod}(\Sigma)$. This is because of the so-called Tits-alternative: a subgroup $G\leqslant\operatorname{PMod}(\Sigma)$ is either virtually Abelian or it contains free groups [30, Theorem A]. In particular, up to taking powers, the subgroup generated by two mapping classes is almost always a free group.
@@
The idea is to show that any non-central $N\triangleleft\operatorname{Mod}(\Sigma_{g}^{b})$ contains pseudo-Anosov maps [^1] See [10, §13.2.3] for a definition., which is known since the mid 1980’s – see [28, Lemma 2.5]. This is also a consequence of the existence of all pseudo-Anosov normal subgroups of $\operatorname{Mod}(\Sigma_{g}^{b})$, a fact first established by Dahmani–Guirardel–Osin [7, Theorem 2.31].
@@ 2.2. The Torelli group & its Subgroups
The subgroup $\operatorname{SIP}(\Sigma)\leqslant\mathcal{I}(\Sigma)$ generated by simple intersection maps remains less well understood than $\mathcal{I}(\Sigma)$ and $\mathcal{K}(\Sigma)$. This is a normal subgroup of $\operatorname{Mod}(\Sigma)$: given a simple intersection pair $(a,b)$ and $f\in\operatorname{Mod}(\Sigma)$,
@@
When $\Sigma=\Sigma_{g}^{1}$, the images of simple intersection maps under the Johnson homomorphism were computed independently by Putman [35], Church [6] and Childers [5, Main Result 1]. Childers also computed the image of $\operatorname{SIP}(\Sigma_{g}^{1})$ under the so-called Birman–Craggs–Johnson homomorphism [5, Main Result 4].
@@ 3. Commutation Relations in the Mapping Class Group
The latter is a Gromov-hyperbolic graph on which $\operatorname{PMod}(\Sigma)$ acts by isometries [29]. Given the disjointness relations from §2.1, $\mathcal{C}(\Sigma)$ and $\Gamma(\operatorname{PMod}(\Sigma))$ are related by means of the embedding $\mathcal{C}(\Sigma)\hookrightarrow\Gamma(\operatorname{PMod}(\Sigma))$ taking $a\subseteq\Sigma$ to its Dehn twist $T_{a}$.
@@
Now each $W_{i+1}/W_{i}$ may be decomposed as a tensor product $W_{i+1}/W_{i}=W_{i,1}\otimes\cdots\otimes W_{i,n}$, where $\rho_{ij}:H_{j}\to\operatorname{GL}(W_{ij})$ is an irreducible representation and $\rho_{i}(h_{1},\ldots,h_{n})=\rho_{i,1}(h_{1})\otimes\cdots\otimes\rho_{i,n}(h_{n})$ – see, for example, [24, Proposition 2.3.23]. In this setting,
$\dim W_{i+1}/W_{i}=(\dim W_{i,1})\cdots(\dim W_{i,n})\geqslant 2^{\#\{j:\dim W_{ij}\geqslant 2\}}\geqslant 2\cdot\#\{j:\dim W_{ij}\geqslant 2\},$
for $2^{k}\geqslant 2k$ for all integers $k\geqslant 0$. In particular, $d=\dim W_{1}/W_{0}+\cdots+\dim W_{p+1}/W_{p}\geqslant 2\cdot\#\{(i,j):\dim W_{ij}\geqslant 2\}$.
If $\dim W_{ij}=1$ then $\rho_{ij}(h_{j})$ is a scalar operator for all $h_{j}\in H_{j}$, so that $\rho_{i}(1,\ldots,h_{j},\ldots,1)=1\otimes\cdots\otimes\rho_{ij}(h_{j})\otimes\cdots\otimes 1$ is also a scalar operator. Assume we can find $j\leqslant n$ with $\dim W_{ij}=1$ for all $i$. Then the matrix $\rho(\pi(1,\ldots,h_{j},\ldots,1))$ is upper triangular with respect to the basis from (3.3) for all $h_{j}\in H_{j}$. Since the group of upper triangular matrices is solvable, it follows $\pi(H_{i}^{(k)})\leqslant\ker\rho$ for some $k$.
But $H_{i}$ is non-solvable and so $\pi(H_{i}^{(k)})\cong H_{i}^{(k)}\neq 1$, contradicting the assumption that $\rho$ is faithful. This means that, for each $j\leqslant n$ we can find $i$ such that $\dim W_{ij}\geqslant 2$. Hence $d\geqslant 2\cdot\#\{(i,j):\dim W_{ij}\geqslant 2\}\geqslant 2n$, as desired. ∎
@@
Figure 4. The curves $a_{1},\ldots,a_{k},b_{1},\ldots,b_{k}\subseteq\mathbb{D}_{n}$.
In that case, consider $a_{1},\ldots,a_{k},b_{1},\ldots,b_{k}\subseteq\mathbb{D}_{n}$ as in Figure 4. It follows from the discussion above that the subgroup generated by $T_{a_{1}},\ldots,T_{a_{k}},T_{b_{1}},\ldots,T_{b_{k}}\in\operatorname{PMod}(\mathbb{D}_{n})$ is a quotient of $G_{0,k}$, the direct product of $k$ copies of $F_{2}$. What is more, the projection of each $F_{2}$-factor onto its image in $\operatorname{PMod}(\Sigma_{0,n}^{1})$ is an isomorphism. Proposition 3.2 then says $d\geqslant 2k=n-1$, as desired. ∎
@@
Let $\rho:\operatorname{Mod}(\Sigma_{g}^{1})\hookrightarrow\operatorname{GL}_{d}(\mathbb{C})$ be a faithful representation and $g^{\prime}=\lfloor\nicefrac{{g}}{{n}}\rfloor$. Take $d_{\operatorname{min}}=\min\{d(E):E\text{ is a cyclic extension of }\operatorname{Mod}(\Sigma_{g^{\prime},1})\}$. We want to show $d\geqslant n\cdot d_{\operatorname{min}}$. By passing to a smaller subsurface $\Sigma_{n\cdot g^{\prime}}^{1}\subseteq\Sigma_{g}^{1}$ if necessary, we may assume $g=n\cdot g^{\prime}$ with $g^{\prime}\geqslant 2$. In that case, we may view $\Sigma_{g}^{1}$ as an $(n+1)$-hole sphere with $n$ copies $S_{1},\ldots,S_{n}\subseteq\Sigma_{g}^{1}$ of $\Sigma_{g^{\prime}}^{1}$ attached along their boundaries.
@@
so that the action $\rho_{i}:\operatorname{Mod}(S_{1})\times\cdots\times\operatorname{Mod}(S_{n})\to\operatorname{GL}(W_{i+1}/W_{i})$ of $\operatorname{Mod}(S_{1})\times\cdots\times\operatorname{Mod}(S_{n})$ on each successive quotient is irreducible. Set $W_{i+1}/W_{i}=W_{i,1}\otimes\cdots\otimes W_{i,n}$, where $\rho_{ij}:\operatorname{Mod}(S_{j})\to\operatorname{GL}(W_{ij})$ is an irreducible representation and $\rho_{i}(f_{1},\ldots,f_{n})=\rho_{i,1}(f_{1})\otimes\cdots\otimes\rho_{i,n}(f_{n})$ [24, Proposition 2.3.23].
@@
$\begin{split}d&=\dim W_{1}/W_{0}+\cdots+\dim W_{p+1}/W_{p}\\ &\geqslant d_{\operatorname{min}}^{\#\{j:\dim W_{1,j}\geqslant d_{\operatorname{min}}\}}+\cdots+d_{\operatorname{min}}^{\#\{j:\dim W_{p,j}\geqslant d_{\operatorname{min}}\}}\\ &\geqslant d_{\operatorname{min}}\cdot\#\{(i,j):\dim W_{ij}\geqslant d_{\operatorname{min}}\}\\ &\geqslant n\cdot d_{\operatorname{min}},\end{split}$
@@
Take $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ with $d$ small enough. In §4 and §5 we will show that, unless $\rho$ kills the generators of $\operatorname{SIP}_{0}(\Sigma)$, the matrices $M_{i}=\rho(T_{a_{i}})-1$ and $N_{j}=\rho(T_{b_{j}})-1$ satisfy the “annihilation relations”
@@
Let $M_{1},\ldots,M_{n},N_{1},\ldots,N_{n}\in M_{d}(\mathbb{C})$ be nonzero operators subject to the annihilation relations (3.5). Then $\dim\sum_{i}(\operatorname{range}M_{i}+\operatorname{range}N_{i})\geqslant 3n-d$. In particular, $2d\geqslant 3n$.
@@ 4. Eigenspaces of $T_{a}$
Let $\Sigma$ be a genus $g$ surface and $\rho:\operatorname{PMod}(\Sigma)\to\operatorname{GL}_{d}(\mathbb{C})$ satisfying either (1) or (2) from Proposition 4.2. Take $a,b\subseteq\Sigma$ disjoint with $a$ nonseparating. Then $(L_{a}-1)(L_{b}-1)=0$.
@@
Now since $\rho$ is nontrivial, $\dim E_{1}^{a}>2g-2$ by Proposition 4.2 and thus $\dim\mathbb{C}^{d}/E^{a}_{1}<2g-2$. It follows from Theorem 2.7 that $\bar{\rho}(f)=1$. Given any $b\subseteq\Sigma$ disjoint from $a$, $b\subseteq\Sigma_{a}$ and so we may write
@@
Proof.
The map $A-\lambda$ induces an injective linear map $(A-\lambda):E_{\lambda,k+1}/E_{\lambda,k}\hookrightarrow E_{\lambda,k}/E_{\lambda,k-1}$ for all $k$. ∎
@@
On the one hand, $\operatorname{Ext}^{1}_{\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]}(H_{\mathbb{Z}},H_{\mathbb{Z}})=H^{1}(\operatorname{Mod}(\Sigma^{\prime});\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},H_{\mathbb{Z}}))$ [4, Proposition 2.2]. Here $\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]$ acts on $\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},H_{\mathbb{Z}})$ by $f\cdot A=\Psi(f)\circ A\circ\Psi(f)^{-1}$ for all $f\in\operatorname{Mod}(\Sigma^{\prime})$ and $A\in\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},H_{\mathbb{Z}})$. Hence $\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},H_{\mathbb{Z}})\cong H_{\mathbb{Z}}^{*}\otimes_{\mathbb{Z}}H_{\mathbb{Z}}$ as $\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]$-modules, where $\mathbb{Z}[\operatorname{Mod}(\Sigma^{\prime})]$ acts on $H_{\mathbb{Z}}^{*}=\operatorname{Hom}_{\mathbb{Z}}(H_{\mathbb{Z}},\mathbb{Z})$ via $f\cdot A=A\circ\Psi(f)^{-1}$.